OpenStudy (anonymous):

what % of 30 of 54? please help! I just need the answer!

OpenStudy (king):

180

OpenStudy (anonymous):

thx

jhonyy9 (jhonyy9):

you can calcule in this way 54 .... 100 % 30 ..... x ------------------- x=(30*100)/54 = 3000/54 ok.???